# Towards Dataset Copyright Evasion Attack against Personalized Text-to-Image Diffusion Models

### v1

@@ Abstract
Text-to-image (T2I) diffusion models have rapidly advanced, enabling high-quality image generation conditioned on textual prompts. However, the growing trend of fine-tuning pre-trained models for personalization raises serious concerns about unauthorized dataset usage. To combat this, dataset ownership verification (DOV) has emerged as a solution, embedding watermarks into the fine-tuning datasets using backdoor techniques. These watermarks remain inactive under benign samples but produce owner-specified outputs when triggered. Despite the promise of DOV for T2I diffusion models, its robustness against copyright evasion attacks (CEA) remains unexplored. In this paper, we explore how attackers can bypass these mechanisms through CEA, allowing models to circumvent watermarks even when trained on watermarked datasets. We first analyze the limitations of potential attacks achieved by backdoor removal, including TPD and T2IShield. In general, TPD, which randomly perturbs textual prompts, lacks consistent effectiveness due to randomness, while T2IShield, which detects watermarked samples via cross-attention differences, fails when watermarks are embedded as local image patches. To overcome these shortcomings, we propose the first copyright evasion attack (i.e., CEAT2I) specifically designed to undermine DOV in T2I diffusion models. Concretely, our CEAT2I comprises three stages: watermarked sample detection, trigger identification, and efficient watermark mitigation. A key insight driving our approach is that T2I models exhibit faster convergence on watermarked samples during the fine-tuning, evident through intermediate feature deviation. Leveraging this, CEAT2I can reliably detect the watermarked samples. Then, we iteratively ablate tokens from the prompts of detected watermarked samples and monitor shifts in intermediate features to pinpoint the exact trigger tokens. Finally, we adopt a closed-form concept erasure method to remove the injected watermark. Extensive experiments show that our CEAT2I effectively evades DOV mechanisms while preserving model performance. The code is available at https://github.com/csyufei/CEAT2I.
Index Terms:
Dataset Ownership Verification, Copyright Evasion Attack, Text-to-Image Diffusion Models.
@@ I Introduction
In recent years, Text-to-image (T2I) diffusion models [15, 43, 45] have made significant progress, revolutionizing the landscape of generative AI. Large pre-trained T2I diffusion models, such as Stable Diffusion [45], have demonstrated impressive capabilities in generating high-quality images from textual prompts. These models have been widely adopted across various domains, from creative industries to scientific visualization, enabling users to produce intricate and realistic images that closely align with provided prompts.
@@
In addition to their remarkable capabilities in generating general images, there is a growing interest in customizing personalized T2I models [19, 13, 46] to produce images in specific themes, such as mimicking a particular artist’s style or replicating a branded visual identity. Personalization is typically achieved by fine-tuning a pre-trained diffusion model using a reference dataset. The result is a personalized/customized model that can generate images with striking fidelity to the desired aesthetic. However, the success of this personalization process heavily relies on access to high-quality fine-tuning datasets. This growing reliance on high-quality datasets has raised serious concerns about unauthorized usage. For example, artists may worry that their work may be used without authorization to fine-tune personalized T2I models, enabling others to generate imitations in their distinctive style, potentially infringing on copyrights and intellectual property. Similarly, organizations that release datasets for limited, non-commercial use (e.g., academic research) are concerned that their data might be misused to fine-tune models for profit. In cases where a suspicious model is found to generate outputs closely resembling a protected dataset, the data owner may suspect misuse but lack conclusive proof, making it difficult to enforce terms of use or pursue legal recourse.
@@
To address this issue, dataset ownership verification (DOV) [31, 32, 65] has emerged as an effective approach to safeguard datasets from the unauthorized use. DOV methods typically employ backdoor-based watermark techniques to embed unique triggers within datasets. It can enable dataset owners to verify whether a suspect model has been trained on the watermarked dataset. Specifically, when T2I diffusion models use the backdoor-based watermarked dataset during the fine-tuning process, they behave normally when access to benign samples. However, when the owner-specified triggers present, they either generate a predefined global image [63, 49, 9], such as a logo, or a local patch within an image [63], such as a signature. These watermarks are designed to leave no observable trace during regular use but activate under owner-specified triggers. By leveraging such techniques, DOV can provide a viable means for dataset owners to assert their dataset ownership and take necessary actions against the unauthorized dataset usage.
@@
Despite the advancements in DOV methods, there remains a lack of approaches to assess their robustness. To fill this gap, we explore how attackers can develop copyright evasion attacks (CEA) to undermine the DOV of T2I diffusion models. Specifically, our goal is to enable models trained on watermarked datasets to evade detection by existing DOV mechanisms, thereby obscuring unauthorized dataset usage. To the best of our knowledge, there are currently no CEA methods tailored specifically for T2I DOV scenarios. However, since DOV approaches often rely on backdoor-based watermark techniques, we begin by analyzing the limitations of current backdoor removal techniques in T2I diffusion models, including textual perturbation defense (TPD) [8] and T2IShield [54]. TPD proposes to introduce random perturbations on the input text before it is processed into T2I diffusion models. The goal is to disrupt the activation of hidden triggers that produce watermarked outputs. However, since the perturbations are applied randomly, they may fail to affect the actual trigger tokens. Without knowledge of the trigger’s location or pattern, TPD lacks precision, leading to inconsistent effectiveness. Consequently, random perturbation alone is often insufficient to reliably suppress the watermark response. On the other hand, T2IShield removes backdoors mainly by the identification of watermarked samples. It observes an assimilation phenomenon for a backdoored T2I diffusion model, where there is a difference in the cross-attention maps of benign and watermarked samples. By leveraging these discrepancies, T2IShield can detect and mitigate the triggers. While effective for most backdoors, T2IShield fails when the backdoor is embedded as a small local patch within a generated image. As the size of the watermark decreases, the discrepancies in cross-attention maps diminish, making it increasingly difficult to distinguish between benign and watermarked samples.
@@
To address the limitations, we propose an effective copyright evasion attack (CEAT2I) tailored for the DOV of T2I diffusion models. CEAT2I is designed to ensure that a watermark-free model can be obtained even when fine-tuning on a watermarked dataset. Our approach is built by three key components: watermarked sample detection, trigger identification, and efficient watermark mitigation. A critical challenge in countering DOV lies in the accurate detection of watermarked samples, which is an aspect inadequately addressed by prior methods, particularly in cases involving subtle local watermarks. To overcome this, CEAT2I introduces a robust detection strategy that works effectively for both global image watermarks and localized patches. Our key observation is that during fine-tuning on a watermarked dataset, T2I diffusion models converge significantly faster to watermarked samples in intermediate representations. Specifically, the $\mathcal{L}_{2}$ loss between the original and fine-tuned models’ feature values is noticeably higher for watermarked samples than benign ones in the early fine-tuning epochs. By analyzing the convergence difference of intermediate features, we can effectively distinguish watermarked samples from benign ones. Subsequently, to locate the trigger within a watermarked sample, we iteratively remove each word from the input texts of recognized watermarked samples while keeping the rest unchanged. We then compare the resulting intermediate features with those generated from the full prompt. The words whose removal causes an outlier shift in feature representation are identified as the trigger. Finally, using the identified triggers and the fine-tuned model, we apply a closed-form concept erasure technique to neutralize their effects. Given a watermarked dataset, CEAT2I enables the recovery of a model that behaves as if it were trained on a benign dataset, without requiring additional fine-tuning. The compared effects of existing attacks and our proposed CEAT2I on DOV in T2I diffusion models are briefly shown in Fig. 1.
@@
  • Based on previous findings, we propose a simple yet effective method (dubbed “CEAT2I”) for T2I diffusion models. Notably, CEAT2I is robust to both global watermarks and local patch watermarks for DOV.
@@ II-A Text-to-Image Diffusion Model
II-B Data Protection
Data protection [12, 6, 21, 31, 40, 23, 28, 38, 33, 66] is a fundamental research area aimed at preventing unauthorized data usage and safeguarding data privacy. Existing protection methods can be broadly categorized into private data protection and public data protection, which depends on the nature of the protected data.
Most traditional methods focus on protecting private data, employing techniques such as encryption, digital watermarking, and differential privacy. Encryption [10, 20, 60] secures data using a secret key, ensuring that only authorized users can access and decrypt the information. Digital watermarking [26, 18, 39, 5] embeds owner-specific patterns into digital assets, allowing verification of ownership by detecting predefined watermarks. Differential privacy [4, 48, 61, 53, 47, 37] introduces noise during the model training to prevent the leakage of sensitive information from gradients or model parameters. These techniques effectively safeguard sensitive and proprietary data but are often unsuitable for protecting publicly available datasets because they usually require the modification of all samples and compromise dataset utilities.
@@ II-B Data Protection
Protecting public data, such as datasets from social media or open-source repositories, is a relatively recent challenge, due to the black-box verification for data owners. Existing solutions fall into two main categories: unlearnable examples [21, 44, 22] and dataset ownership verification (DOV) [31, 32, 50, 55, 27]. Unlearnable examples poison the dataset by altering all samples in a way that prevents machine learning models from learning meaningful representations. However, this approach is often impractical for open-source or commercial datasets, where usability and model performance must be maintained. Dataset ownership verification (DOV) provides a more practical solution by embedding identifiable patterns into datasets to verify whether a suspicious third-party model has been trained on the protected data. DOV typically involves two key stages: dataset watermarking and ownership verification. The most commonly used technique for dataset watermarking is backdoor-based watermarking, where models trained on watermarked datasets exhibit distinct behaviors (e.g., misclassification or generation of predefined patterns) when presented with specific triggers while performing normally on benign samples. Consequently, one of the critical challenges in DOV lies in designing effective and robust backdoor-based watermarks that remain detectable while minimizing their impact on standard model functionality.
II-C Dataset Ownership Verification on T2I Diffusion Models
Dataset ownership verification (DOV) [31, 32, 50, 17, 55, 7, 27, 16] typically adopts backdoor-based watermark techniques to protect training datasets from unauthorized use. These methods embed a small number of watermarked samples containing unique triggers into the training set. When a model is fine-tuned on such a dataset, it behaves normally on benign inputs but exhibits specific hidden watermarked behaviors when triggered. Most existing DOV approaches have been primarily developed for image classification datasets [27, 16, 17], where the watermarked behavior typically involves predicting a target label when the trigger is present. Differently, when applied to T2I diffusion models, these DOV methods typically aim to manipulate the model into generating either a specific local patch within an image [63] or a global target image [63, 49, 9, 65, 52, 11] when given an input containing the trigger. Rickrolling [49] first demonstrated that visually similar non-Latin characters (homoglyphs) could serve as triggers to generate a target image from an unrelated prompt. BadT2I [63] applies full model fine-tuning to achieve localized or full-image manipulation. VillanDiffusion [9] proposes to fine-tune the U-Net component of diffusion models to enable a flexible and unified framework compatible with different samplers and text triggers. These techniques effectively establish an association between a trigger and either a specific local patch (e.g., a signature) or an entire target image (e.g., a logo). Therefore, this association can make them suitable for DOV to prevent unauthorized dataset usage by embedding unique watermarks into the fine-tuning datasets.
Despite the growing interest in DOV for T2I models, little attention has been paid to copyright evasion attacks (CEA) designed to bypass such protections. Since DOV relies heavily on backdoor-based watermarks, we begin by analyzing the limitations of existing backdoor removal strategies, including Textual Perturbation Defense (TPD) [8] and T2IShield [54]. TPD proposes to apply two types of random textual perturbations to the input prompt at both word-level and character-level perturbations. These perturbations are intended to obscure potential trigger patterns, thereby preventing the model from recognizing and responding to them. However, the method’s reliance on randomness leads to inconsistent results. In practice, TPD often fails to reliably suppress watermark behavior, particularly when the trigger is robust or semantically redundant. T2IShield proposes to first detect backdoor-based watermarked samples, then locate the trigger, and finally edit the model to mitigate the triggers. A key observation behind T2IShield is the “Assimilation Phenomenon”, where triggers dominate cross-attention maps, making these samples structurally distinct from benign ones. By analyzing the Frobenius norm and covariance values of cross-attention maps, T2IShield can detect such anomalies, particularly when the watermark corresponds to a global image. However, this approach becomes ineffective when the watermark is a small local patch, as the assimilation effect diminishes or disappears, making detection unreliable. Besides, the trigger localization in T2IShield relies on additional models, such as CLIP [42] and DinoV2 [36]. Given the limitations of current backdoor removal techniques, there is currently no effective CEA approach for DOV, highlighting the need for an effective method to counteract DOV mechanisms in T2I diffusion models.
@@ IV-C Watermarked Sample Detection
Watermarked samples are the foundation of backdoor-based watermark injection in T2I diffusion models, as they can enable the specific trigger-target associations embedded into the model during fine-tuning. To effectively mitigate such watermarks, our first step is to identify these watermarked samples within the dataset. Our detection approach is based on a key empirical observation: watermarked samples exhibit distinct learning dynamics compared to benign ones. When a model is fine-tuned on a dataset containing backdoor-based watermarks, the presence of the trigger-target correlations causes the model to adapt its internal representations more rapidly for watermarked samples. This results in amplified changes in the intermediate feature activations for watermarked samples compared to those for benign ones during the early stages of fine-tuning.
Let $f_{\bm{\theta}}^{i}(\bm{z}_{t},t,\bm{c})$ and $f_{\bm{\theta_{w}}}^{i}(\bm{z}_{t},t,\bm{c})$ denote the feature activations at the $i$-th layer of the original and fine-tuned T2I diffusion models at an early epoch $T_{e}$, respectively. For a given image-text pair $(\bm{x},y)$ and a diffusion timestep $t$, we compute the feature deviation at layer $i$ using the $\mathcal{L}_{2}$ distance:
@@
$(\bm{x},y)=\begin{cases}(\bm{x}_{w},y_{w})\in\mathcal{D}_{w}&\text{if }\sum_{i=1}^{N}\mathbf{1}\{\mathcal{L}_{f}^{i}>\alpha_{1}\}>\alpha_{2},\\ (\bm{x}_{b},y_{b})\in\mathcal{D}_{b}&\text{otherwise},\end{cases}$ (4)

### v2

@@ Abstract
Text-to-image (T2I) diffusion models enable high-quality image generation conditioned on textual prompts. However, fine-tuning these pre-trained models for personalization raises concerns about unauthorized dataset usage. To address this issue, dataset ownership verification (DOV) has recently been proposed, which embeds watermarks into fine-tuning datasets via backdoor techniques. These watermarks remain dormant on benign samples but produce owner-specified outputs when triggered. Despite its promise, the robustness of DOV against copyright evasion attacks (CEA) remains unexplored. In this paper, we investigate how adversaries can circumvent these mechanisms, enabling models trained on watermarked datasets to bypass ownership verification. We begin by analyzing the limitations of potential attacks achieved by backdoor removal, including TPD and T2IShield. In practice, TPD suffers from inconsistent effectiveness due to randomness, while T2IShield fails when watermarks are embedded as local image patches. To this end, we introduce CEAT2I, the first CEA specifically targeting DOV in T2I diffusion models. CEAT2I consists of three stages: (1) motivated by the observation that T2I models converge faster on watermarked samples with respect to intermediate features rather than training loss, we reliably detect watermarked samples; (2) we iteratively ablate tokens from the prompts of detected samples and monitor feature shifts to identify trigger tokens; and (3) we apply a closed-form concept erasure method to remove the injected watermarks. Extensive experiments demonstrate that CEAT2I effectively evades state-of-the-art DOV mechanisms while preserving model performance. The code is available at https://github.com/csyufei/CEAT2I.
@@ I Introduction
In recent years, Text-to-image (T2I) diffusion models [13, 49, 51] have made significant progress. Large pre-trained T2I diffusion models, such as Stable Diffusion [51], have demonstrated impressive capabilities in generating high-quality images from textual prompts. These models have been widely adopted across various domains, from creative industries to scientific visualization.
@@
In addition to their remarkable capabilities in generating general images, there is a growing interest in customizing personalized T2I models [19, 10, 52] to produce images in specific themes, such as mimicking a particular artist’s style. Personalization is typically achieved by fine-tuning a pre-trained diffusion model using a reference dataset. The result is a customized model that can generate images with striking fidelity to the desired aesthetic. However, the success of this personalization process heavily relies on access to high-quality fine-tuning datasets. This growing reliance on high-quality datasets has raised serious concerns about unauthorized usage. For example, artists may worry that their work may be used without authorization to fine-tune personalized T2I models, enabling others to generate imitations in their distinctive style. Similarly, organizations that release datasets for limited, non-commercial use (e.g., academic research) are concerned that their data might be misused to fine-tune models for profit. In cases where a suspicious model is found to generate outputs closely resembling a protected dataset, the data owner may suspect misuse but lack conclusive proof, making it difficult to enforce terms of use or pursue legal recourse.
@@
To address this issue, dataset ownership verification (DOV) [34, 37, 72, 36] has emerged as an effective approach to safeguard datasets from the unauthorized use. DOV methods typically employ backdoor-based watermark techniques to embed unique triggers within datasets. It can enable dataset owners to verify whether a suspect model has been trained on the watermarked dataset. Specifically, when T2I diffusion models use the backdoor-based watermarked dataset during the fine-tuning process, they behave normally when access to benign samples. However, when the owner-specified triggers present, they either generate a predefined global image [70, 55, 6], such as a logo, or a local patch within an image [70], such as a signature. These watermarks are designed to leave no observable trace during regular use but activate under owner-specified triggers. By leveraging such techniques, DOV can provide a viable means for dataset owners to assert their dataset ownership and take necessary actions against the unauthorized dataset usage.
@@
Despite recent progress in DOV methods, their robustness has largely been evaluated only against naive strategies (e.g., fine-tuning), with no practical method to assess their resilience against more sophisticated and adaptive adversaries in real-world. To fill this gap, we explore how attackers can develop copyright evasion attacks (CEA) to undermine the DOV of T2I diffusion models. Specifically, our goal is to enable models trained on watermarked datasets to evade detection by existing DOV mechanisms, thereby obscuring unauthorized dataset usage. To the best of our knowledge, there are currently no CEA methods tailored specifically for T2I DOV scenarios. However, since DOV approaches often rely on backdoor-based watermark techniques, we begin by analyzing the limitations of current backdoor removal techniques in T2I diffusion models, including textual perturbation defense (TPD) [5] and T2IShield [61]. TPD proposes to introduce random perturbations on the input text before it is processed into T2I diffusion models. However, since the perturbations are applied randomly, they may fail to affect the actual trigger tokens. Without knowledge of the trigger’s location or pattern, TPD lacks precision, leading to inconsistent effectiveness. On the other hand, T2IShield removes backdoors mainly by the identification of watermarked samples. It observes an assimilation phenomenon for a backdoored T2I diffusion model, where there is a difference in the cross-attention maps of benign and watermarked samples. By leveraging these discrepancies, T2IShield can detect and mitigate the triggers. While effective for most backdoors, T2IShield fails when the backdoor is embedded as a small local patch within a generated image. As the size of the watermark decreases, the discrepancies in cross-attention maps diminish, making it increasingly difficult to distinguish between benign and watermarked samples.
@@
To overcome the aforementioned limitations, we propose CEAT2I, an effective copyright evasion attack tailored for DOV in T2I diffusion models. CEAT2I is specifically designed to obtain a watermark-free model even when fine-tuned on watermarked datasets. It consists of three key components: watermarked sample detection, trigger identification, and efficient watermark mitigation. A critical challenge in undermining DOV is the accurate detection of watermarked samples, which prior methods fail to address, especially for subtle local watermarks. CEAT2I introduces a robust detection strategy that is effective against both global watermarks and localized patches. The key insight is that, during fine-tuning, T2I diffusion models converge significantly faster on watermarked samples in intermediate representations, rather than in training loss. In particular, the $\mathcal{L}_{2}$ distance between feature values of the original and fine-tuned models is consistently larger for watermarked samples than benign ones during early epochs. By leveraging this convergence disparity, CEAT2I can reliably distinguish watermarked samples. Once identified, the corresponding triggers are located via feature deviations by iteratively ablating words from the input prompts of detected samples while keeping the remaining text unchanged. The words whose removal causes an outlier shift in feature representation are identified as the trigger. Finally, given the detected triggers and the fine-tuned model, we employ a closed-form concept erasure method to neutralize their effects. A comparison of existing attacks and our proposed CEAT2I on DOV in T2I diffusion models is illustrated in Fig. 1.
@@
  • Building on these findings, we propose a simple yet effective method, i.e., CEAT2I, for T2I diffusion models. CEAT2I demonstrates robustness against both global and local patch watermarks in DOV, primarily due to the effectiveness of its watermarked sample detection.
@@ II-A Text-to-Image Diffusion Model
II-B Dataset Ownership Verification
@@
Data protection [2, 8, 31, 35, 39] aims to prevent unauthorized data usage and safeguard data privacy. Existing approaches are generally divided into private and public data protection. Private data protection, such as encryption [7, 20, 68], digital watermarking [29, 17, 44], and differential privacy [1, 73, 43], secure sensitive information by restricting access, embedding ownership marks, or adding noise to prevent leakage. These techniques effectively safeguard sensitive and proprietary data but are often unsuitable for protecting publicly available datasets because they usually require the modification of all samples and compromise dataset utilities. Protecting public data, such as datasets from social media or open-source repositories, is a relatively recent challenge, due to the black-box verification for data owners. Existing solutions fall into two main categories: unlearnable examples and dataset ownership verification. Unlearnable examples [21, 50, 24] poison the dataset by altering all samples in a way that prevents machine learning models from learning meaningful representations. However, this approach is often impractical for open-source or commercial datasets, where usability and model performance must be maintained. Dataset ownership verification (DOV) [34, 37, 62, 4] provides a more practical solution by embedding identifiable patterns into datasets to verify whether a suspicious third-party model has been trained on the protected data. DOV typically adopts backdoor-based watermark techniques to protect training datasets from unauthorized use. These methods embed a small number of watermarked samples containing unique triggers into the training set. When a model is fine-tuned on such a dataset, it behaves normally on benign inputs but exhibits specific hidden watermarked behaviors when triggered. In particular, unlike malicious backdoor attacks whose purpose is to induce unsafe model behaviors [14], the goal of DOV is to enable data owners to prove the presence of their data in unauthorized model training based on the distinctive inference behaviors (e.g., backdoor) on defender-specified verification samples. Moreover, backdoor attacks may manipulate the entire training pipeline (e.g., loss functions and learning schedules), whereas DOV is strictly limited to modifying only the watermarked dataset supplied by the data owner.
@@
Most existing DOV approaches have been primarily developed for image classification datasets [54, 15, 16, 46], where the watermarked behavior typically involves predicting a target label when the trigger is present. Differently, when applied to T2I diffusion models, these DOV methods typically aim to manipulate the model into generating either a specific local patch within an image [70] or a global target image [72, 60, 9] when given an input containing the trigger. Rickrolling [55] first demonstrated that visually similar non-Latin characters (homoglyphs) could serve as triggers to generate a target image from an unrelated prompt. BadT2I [70] applies full model fine-tuning to achieve localized or full-image manipulation. VillanDiffusion [6] proposes to fine-tune the U-Net component of diffusion models to enable a flexible and unified framework compatible with different samplers and text triggers. These techniques effectively establish an association between a trigger and either a specific local patch (e.g., a signature) or an entire target image (e.g., a logo). Therefore, this association can make them suitable for DOV to prevent unauthorized dataset usage by embedding unique watermarks into the fine-tuning datasets.
Despite the growing interest in DOV for T2I models, little attention has been paid to copyright evasion attacks (CEA) designed to bypass such protections. Since DOV relies heavily on backdoor-based watermarks, we begin by analyzing the limitations of existing backdoor removal strategies, including Textual Perturbation Defense (TPD) [5] and T2IShield [61]. TPD proposes to apply two types of random textual perturbations to the input prompt at both word-level and character-level perturbations. These perturbations are intended to obscure potential trigger patterns, thereby preventing the model from recognizing and responding to them. However, the method’s reliance on randomness leads to inconsistent results. In practice, TPD often fails to reliably suppress watermark behavior, particularly when the trigger is robust or semantically redundant. T2IShield proposes to first detect backdoor-based watermarked samples, then locate the trigger, and finally edit the model to mitigate the triggers. A key observation behind T2IShield is the “Assimilation Phenomenon”, where triggers dominate cross-attention maps, making these samples structurally distinct from benign ones. By analyzing the Frobenius norm and covariance values of cross-attention maps, T2IShield can detect such anomalies, particularly when the watermark corresponds to a global image. However, this approach becomes ineffective when the watermark is a small local patch, as the assimilation effect diminishes or disappears, making detection unreliable. Besides, the trigger localization in T2IShield relies on additional models, such as CLIP [48] and DinoV2 [42]. Given the limitations of current backdoor removal techniques, there is currently no effective CEA [12, 54] for T2I models, highlighting the need for an effective method to counteract DOV mechanisms in T2I models.
@@ IV-C Watermarked Sample Detection
In the first stage, CEAT2I aims to identify watermarked samples within the training dataset. Specifically, compared to benign samples, watermarked samples exhibit faster feature convergence during fine-tuning. Therefore, CEAT2I classifies those with larger feature shifts as watermarked.
Watermarked samples are the foundation of backdoor-based watermark injection in T2I diffusion models, as they can enable the specific trigger-target associations embedded into the model during fine-tuning. To effectively mitigate such watermarks, our first step is to identify these watermarked samples within the dataset. Inspired by existing backdoor removal techniques, such as ABL [32], our approach builds on a key empirical observation: watermarked samples exhibit distinct learning dynamics compared to benign ones. ABL relies on loss-based detection but the highly smooth loss landscapes of T2I diffusion models [66, 18] diminish the discriminative power of loss-based separation. To address this problem, we introduce a feature-based detection method specifically for T2I diffusion models. From the perspective of information bottleneck (IB) [57, 58], fine-tuning encourages each feature activation $Z$ to preserve only those aspects of the input text $Y$ that are relevant for generating the output image $X$. Formally, the IB objective seeks to minimize $I(Z;Y)-\beta\cdot I(Z;X)$, where training proceeds by enhancing the relevance for generation $I(Z;X)$ while discarding superfluous input information $I(Z;Y)$. Watermarked samples [56] usually embed highly discriminative and low-entropy label information, rendering the generation of $X$ nearly deterministic, i.e., low conditional entropy $H(X|Y)$. Consequently, such samples require lower representational relevance $I(Z;Y)$ to achieve the target predictive mutual information $I(Z;X)$ compared to benign samples. During fine-tuning, these high-gain directions are therefore preferentially amplified: the mutual information $I(Z;X)$ for watermarked samples increases rapidly, whereas benign samples demand more extensive fine-tuning.
As a result, when a model is fine-tuned on a dataset containing backdoor-based watermarks, the presence of the trigger-target correlations causes the model to adapt its internal representations more rapidly for watermarked samples. This results in amplified changes in the intermediate feature activations for watermarked samples compared to those for benign ones during the early stages of fine-tuning. Let $f_{\bm{\theta}}^{i}(\bm{z}_{t},t,\bm{c})$ and $f_{\bm{\theta_{w}}}^{i}(\bm{z}_{t},t,\bm{c})$ denote the feature activations at the $i$-th layer of the original and fine-tuned T2I diffusion models at an early epoch $T_{e}$, respectively. For a given image-text pair $(\bm{x},y)$ and a diffusion timestep $t$, we compute the feature deviation at layer $i$ using the $\mathcal{L}_{2}$ distance:
@@
[…]